Related to Checkpoint 11.1 and Checkpoint 11.4) (NPV and IRR calculation) East Coast Television is considering a project with an initial outlay of SX (you will ave to determine this amount). It is ex ppropriate discount rate for this project is 11 percent. If the project has an internal rate of return of 13 percent, what is the project's net present value? that the project will produce a positive cash flow of $48,000 a year at the end of each year for the next 14 years. The . If the project has an internal rate of return of 13% then the pro eers initial ou ay is S (Round to the nearest cent.
Help the internal rate of return on the following project An initial outlay of $9,000 resulting in a cash inflow of $1,700 at the end of year 1, $4.900 at the end of year 2, and $8,400 at the end of year 3 This project's internal rate of return is[ % (Round to two decimal places)
Answer #1
1) The project's initial outlay = PV of the cash inflows discounted at IRR = 48000*(1.13^14-1)/(0.13*1.13^14) = $     3,02,519.43
NPV = -302519.43+48000*(1.11^14-1)/(0.11*1.11^14) = $         32,610.10
2) IRR is that discount rate for which NPV = 0. Such a discount rate is to be found out by trial
and error.
Year Cash flow PVIF at 25% PV at 25% PVIF at 24% PV at 24% PVIF at 23% PV at 23%
0 -9000 1 $    -9,000.00 1.00000 $ -9,000.00 1.00000 $ -9,000.00
1 1700 0.80000 $      1,360.00 0.80645 $   1,370.97 0.81301 $    1,382.11
2 4900 0.64000 $      3,136.00 0.65036 $   3,186.78 0.66098 $    3,238.81
3 8400 0.51200 $      4,300.80 0.52449 $   4,405.69 0.53738 $    4,514.02
$        -203.20 $       -36.55 $        134.95
IRR = 23%+1%*134.95/(134.95+36.55) = 23.79%